OpenStudy (anonymous):

PLEASE HELP c:

OpenStudy (anonymous):

Mark, Jessica, and Nate each downloaded music from the same website. Mark downloaded 10 songs in total consisting of pop, rock, and hip hop. Jessica downloaded five times as many pop songs, twice as many rock, and three times as many hip hop songs as Mark. She downloaded 28 songs total. If Nate downloaded 20 songs total with three times as many pop songs, three times as many rock songs, and the same number of hip hop songs as Mark, how many songs of each type did Mark download? Mark downloaded 2 pop songs, 5 rock songs, and 3 hip hop songs. Mark downloaded 3 pop songs, 2 rock songs, and 5 hip hop songs. Mark downloaded 4 pop songs, 2 rock songs, and 4 hip hop songs. Mark downloaded 1 pop songs, 4 rock songs, and 5 hip hop songs.

OpenStudy (anonymous):

@Jack1

OpenStudy (uri):

Google thinks...the same :o

OpenStudy (jack1):

Mark: x + y + z =10 (pop+ rock + hip hop) Jess: 5x + 2y + 3z =28 Nate: 3x + 3y + z =20

OpenStudy (jack1):

so again, substitution gives u with x= 1, y= 4, z= 5

OpenStudy (anonymous):

Thank you. I like being explained how to do it so I cant continue to do it and not just given the answer.

OpenStudy (jack1):

all good, same process, it just takes more steps with 3 simultaneous equations

OpenStudy (jack1):

x + y + z =10, EQN 1 5x + 2y + 3z =28, EQN 2 3x + 3y + z =20 EQN 3 so from 1: x = 10 - y - z EQN 1a sub 1a into 2 5x + 2y + 3z =28 5(10 - y - z) + 2y + 3z =28 50 - 5y - 5z +2y + 3z = 28